# AdvQuNN: A Methodology for Analyzing the Adversarial Robustness of Quanvolutional Neural Networks

### v1

@@ Abstract.
Quantum machine learning, Quanvolutional Neural Networks, Quantum Computing, Deep Neural Networks, Convolutional Neural Networks, Adversarial Attacks, Adversarial Robustness.
1. Introduction
Driven by the emergent trends of quantum computing and the massive learning capabilities of Deep Neural Networks (DNNs), Hybrid Quantum Neural Networks (HQNNs) have emerged as high-accurate models to conduct machine learning tasks on currently available quantum devices (Zaman et al., 2023; Markidis, 2023). Due to the reliability constraints of the current existing quantum circuits that limit their sizes, HQNNs combine resource-constrained quantum circuits in quantum layers with classical layers implemented on quantum computers. A common HQNN model is represented by Quanvolutional Neural Networks (QuNNs) (Henderson et al., 2020) that implement a quanvolutional layer followed by classical layers.
Before employing QuNNs in safety-critical systems, it is important to analyze their robustness to different vulnerability threats. While the security of classical DNNs has been extensively studied in recent years (Guesmi et al., 2023), the security of QuNNs has not been analyzed systematically. The recent work in (Sooksatra et al., 2021) has compared QuNNs with analogous classical networks, but did not account for a range of perturbation strength values, while we believe that a more thorough comparison necessitates evaluating the robustness of the models across a spectrum of increased perturbation strengths. Moreover, the work in (Huang et al., 2023) analyzed the effect of different sizes of the kernel and different numbers of filters in QuNNs, which is not sufficient as it is a classical-inspired permutation, and does not fully encapsulate the quantum-specific dynamics, while in our work we investigate the impact of the type of quantum circuit (Ansatz), offering a more inherent quantum perspective.
@@ 1. Introduction
  • We design the AdvQuNN methodology to test different types of adversarial attacks on QuNNs and classical DNN models, varying the quantum circuit type. (Section 3)
  • We compare the robustness of QuNNs to classical CNNs against various types of adversarial attacks. (Section 4.2)
  • We evaluate the impact of the Ansatz architecture on the QuNN robustness (Section 4.3)
Figure 1. Overview of our novel contributions in this work.
2. Background and Related Work
2.1. Classical Convolutional Neural Networks
Convolutional Neural Networks (CNNs) (O’Shea and Nash, 2015; Alzubaidi et al., 2021), particularly effective in visual tasks (Li et al., 2021) like image classification (Rawat and Wang, 2017), object detection (Zhiqiang and Jun, 2017), and semantic segmentation (Guo et al., 2018), rely on convolutional layers to extract spatial features. These layers apply filters to input images, producing feature maps that represent detected features. The convolution operation is mathematically defined as $F(i,j)=(X*K)(i,j)=\sum_{m}\sum_{n}X(m,n)K(i-m,j-n)$, where $F$ is the feature map, $X$ is the input image, and $K$ is the kernel. Subsequent layers combine these features into more complex patterns, following $F^{l}=\sigma\left(\sum_{k=1}^{K}F_{k}^{l-1}*K_{k}^{l}+b_{k}^{l}\right)$, with $\sigma$ as the activation function.
2.2. Quanvolutional Neural Networks
Quanvolutional Neural Networks (QuNNs) (Henderson et al., 2020) represent an intersection between the domains of quantum computing and deep learning that provides an advanced approach to process and understand data. QuNN models substitute the conventional convolutional filter in the CNN with a quantum filter known as the quanvolutional filter. Like classical convolutional layers, the quanvolutional layer applies an operation analogous to an $m\times m$ filter, but it does so by executing an $m^{2}$-qubit quantum circuit (if we use direct pixel encoding as in this work) to produce a downsampled output image. The quanvolutional layer includes three components: data encoding, a parameterized unitary $U(\theta)$, and measurements/decoder $M$.
Data encoding. The input data to the quantum layer, consisting of $n$ features, is encoded into an $n$-qubit Hilbert space using quantum data encoding techniques (LaRose and Coyle, 2020). In this work, we used the angle encoding method where each classical data feature maps to a rotation angle in the quantum circuit, preserving data structure in a less complex quantum circuit. The quantum state of an angle encoded input data $x=[x_{0},x_{1},\dots,x_{n}]^{\intercal}\in\mathbb{R}^{n}$ is represented as
(1) $|\psi_{x}\rangle=\bigotimes_{i}^{n}R\left(\mathbf{\phi}_{i}\right)\left|0^{n}\right\rangle$
@@ 2.2. Quanvolutional Neural Networks
Parameterized unitary $U(\theta)$. A unitary operator $U$, represented by a $2^{n}\times 2^{n}$ matrix, configures an $n$-qubit quantum circuit using single-qubit gates (e.g., ${R_{x},R_{y},R_{z},H,\dots}$) and two-qubit gates (e.g., ${CNOT,CR_{x},CR_{y},CR_{z},\dots}$). These gates enable the creation of various unitary operations $U_{i}$, each forming a distinct quanvolutional filter with specific gate sequences and entanglement patterns. Parameters for these filters can be iteratively updated with classical layer parameters or initialized randomly. In our study, we opt for random initialization to reduce computational demands and simplify the quantum circuit optimization process. The selected parametric circuit $U(\theta)$ is applied to the initial state as follows:
(2) $|\psi_{x}(\theta)\rangle=U(\theta)|\psi_{x}\rangle$
@@
(3) $\langle M\rangle=\langle\psi_{x}(\theta)|M|\psi_{x}(\theta)\rangle=\frac{1}{S}\sum_{i}p_{i}\lambda_{i}$
@@
Figure 2 shows an example of a QuNN architecture applied to the MNIST dataset. A $2\times 2$ square of our input image is first encoded with $R_{y}$ gates into the $2\times 2$ quanvolutional kernel of $4$ qubits. This data then undergoes transformation by a unitary $U(\theta)$, constructed from both parameterized and entangling gates, which may be systematically designed or randomly assembled, as suggested in (Henderson et al., 2020). Finally, we measure the expectation value of the Pauli-$Z$ observable for each qubit. Similarly to the classical convolutional layer, each expectation value contributes to a separate channel in the resulting pixel. Therefore, with four qubits, we obtain four channels. These channels are subsequently flattened and input into a classical dense layer, which determines the likelihood of each digit.
@@
Figure 2. Quanvolutional Neural Network architecture. A 2 by 2 block of the input image of digit $2$ is encoded through a of rotation gates $R_{y}$, producing the state $|\psi_{x}\rangle$. This state is quanvolved with a parameterized unitary gate $U(\theta)$. Subsequently, we measure the state $|\psi_{x}(\theta)\rangle$ and compute the expectation value across all qubits, and assign values to four channels, forming a feature map. Finally, a sequence of fully connected classical layers then processes this map for classification.
2.3. Adversarial Attacks
@@
In Quantum Machine Learning (QML), particularly with quanvolutional neural networks, the issue of adversarial attacks presents a dual challenge of vulnerability and potential robustness. Works by Liu and Wittek (Liu and Wittek, 2020) and Lu et al. (Lu et al., 2020) reveal that these networks, whether they process classical or quantum input data, share vulnerabilities with classical networks, while also presenting a trade-off between algorithmic security and quantum advantages. Conversely, studies by Guan et al. (Guan et al., 2021) and West et al. (West et al., 2023) highlight an intrinsic robustness of quantum models to such attacks. Sooksatra et al. (Sooksatra et al., 2021) and Suryotrisongko et al. (Suryotrisongko et al., 2022) further confirm the superior performance of QuNNs over classical networks in terms of accuracy and resilience against adversarial attacks. This dual aspect of susceptibility and resistance creates an intriguing avenue for developing innovative defense strategies for QML. Our study focuses on investigating a novel approach to evaluate and enhance the robustness of QuNNs.
3. AdvQuNN Methodology
@@
In our methodology, we study the adaptability of QuNNs, focusing on the customizable nature of the unitary operator $U(\theta)$ part of the quanvolutional filter. This flexibility is key in adapting the parametrized quantum circuit, we refer to it interchangeably as an Ansatz, to specific data and learning goals, allowing precise control over the feature extraction process. Such versatility enabled us to propose AdvQuNN, an analysis and design methodology aimed at evaluating the robustness of QuNNs, paving the way for enhanced methods to robustify QuNNs against various adversarial attacks. Our approach involves fine-tuning components such as gate selection and entanglement settings. The following details the specific parametrized circuits used in our study, explaining their design and functionality.
@@
While prior research primarily employed the random Ansatz architecture to evaluate the robustness of QuNN models against adversarial attacks, we decided to assess QuNN models on various Ansatz architectures. We found out that the architecture of the Ansatz affects the expressibility and effectiveness of the quantum circuit and can have a direct impact on the feature extraction capability of QuNNs, hence its robustness. The expressibility of the Ansatz is affected by its rotation gates and entanglement configuration. As quantum entanglement has an important role in the decision boundary formation of the quanvolved data, we explored various entanglement patterns and identified the five most robust architectures for our dataset, see Figure 3. The no_entanglement architecture (Figure 3a) is an Ansatz that does not contain any entanglement and the random architecture (Figure 3e) is a random architecture as used in (Henderson et al., 2020). For the last remaining architectures (Figures 3b, 3c, and 3d), we designed their entanglement pattern to follow the popular nearest-neighbor and all-to-all topologies, but inspired by data encoding technique proved effective for data separation in Quantum Support Vector Machine (QSVM) (Rebentrost et al., 2014), specifically the $ZZ$ feature map that employs $ZZ$ entanglement.
@@
The $ZZ$ entanglement pattern in quantum computing involves a controlled phase shift $e^{-i\theta}$ applied between two qubits when both are in the state $|1\rangle$, leaving other basis states unaffected. This entanglement is created using a unitary transformation $U_{ZZ}(\theta)=$ $e^{-i\theta Z\otimes Z}$, with $Z$ being the Pauli-$Z$ matrix and $\theta$ representing the angle of rotation, and implemented as depicted in red box Figure 3b.
@@
C) $ZZ$ linear entanglement architecture Figure 3 c). In this architecture, each qubit is entangled linearly with its nearest neighbor, forming a chain of entangled qubits with just $n+1$ operations for $n$ qubits. Its corresponding unitary operator is:
(5) $U_{zzlinear}=U^{\mathrm{linear-ent}}\left(\bigotimes_{0}^{n}Rot\left(a,b,c\right)\right)\vspace{-10pt}$
(6) $U^{\text{linear-ent}}:=\prod_{q=1}^{n-1}\mathbf{E}_{q,q+1}\;\;\;\text{where}\;\;\;\mathbf{E}_{q,q+1}=e^{-i\theta Z_{q}Z_{q+1}}$
@@
(7) $U_{zzfull}=U^{\mathrm{full-ent}}\left(\bigotimes_{0}^{n}Rot\left(a,b,c\right)\right)$
(8) $U^{\text{full-ent }}:=\prod_{q=1}^{n-1}\prod_{k=q+1}^{n}\mathbf{E}_{q,k}\;\;\;\text{where}\;\;\;\mathbf{E}_{q,k}=e^{-i\theta Z_{q}Z_{k}}$
@@
(9) $U_{zzstar}=U^{\mathrm{linear-ent}}\left(\bigotimes_{0}^{n}Rot\left(a,b,c\right)\right)\vspace{-10pt}$
(10) $U^{\text{star-ent}}:=\prod_{q=1}^{n-1}\mathbf{E}_{1,q+1}\;\;\;\text{where}\;\;\;\mathbf{E}_{1,q+1}=e^{-i\theta Z_{1}Z_{q+1}}\vspace{-10pt}$
Figure 3. A set of Ansatz architectures considered in the study. Each is labeled as (a) No entanglement, (b) ZZ full entanglement, (c) ZZ linear entanglement, (d) ZZ star entanglement, and (e) Random architecture. The $Rot(a,b,c)$ is a set of three rotation gates $R_{z}(a)R_{y}(b)R_{z}(c)$. The gates $R_{x}$, $R_{y}$, and $R_{z}$ are parameterized.
4. Results and Discussion
4.1. Experimental Setup

### v2

@@ Abstract
Index Terms:
Quantum machine learning, Quanvolutional Neural Networks, Quantum Computing, Deep Neural Networks, Convolutional Neural Networks, Adversarial Attacks, Adversarial Robustness
I Introduction
Driven by the emergent trends of quantum computing and the massive learning capabilities of Deep Neural Networks (DNNs), Hybrid Quantum Neural Networks (HQNNs) have emerged as high-accurate models to conduct machine learning tasks on currently available quantum devices [1, 2]. Due to the reliability constraints of the current existing quantum circuits that limit their sizes, HQNNs combine resource-constrained quantum circuits in quantum layers with classical layers implemented on quantum and classical computers. A common HQNN model is the Quanvolutional Neural Networks (QuNNs) [3] that implement a quanvolutional layer followed by classical layers.
I-A Limitations of State-Of-The-Art and Main Contributions
Before employing QuNNs in safety-critical systems, it is important to analyze their robustness to different vulnerability threats. While the security of classical DNNs has been extensively studied [4], the security of QuNNs has not been analyzed systematically. The recent work in [5] has compared QuNNs with analogous classical networks and the work in [6] studied the impact of quantum hyperparameters on the accuracy of QuNNs. However, they did not account for a range of perturbation strength values, while we believe that a more thorough comparison necessitates evaluating the robustness of the models across a spectrum of increased perturbation strengths. Moreover, the work in [7] analyzed the effect of different sizes of the kernel and different numbers of filters in QuNNs, which is not sufficient as it is a classical-inspired permutation, and does not fully encapsulate the quantum-specific dynamics, while in our work we investigate the impact of the type of quantum circuit (Ansatz), offering a more inherent quantum perspective.
The works in [8] and [9] identify significant adversarial vulnerabilities in conventional neural networks and quantum neural networks, noting that while quantum models exhibit higher resilience and intrinsic robustness due to their architectural properties, they still face considerable adversarial threats. However, neither study delves into the specific design and optimization of quantum circuit architectures to evaluate and enhance robustness. Our work directly addresses this gap by presenting an empirical analysis of various Ansatz architectures in QuNNs. Additionally, the works in [10] and [11] demonstrate that introducing noise, such as depolarization and random rotation noise, enhances the robustness of quantum classifiers against adversarial attacks while satisfying quantum differential privacy. In contrast, our work studied the impact of entanglement in the quantum circuit architecture of QuNNs where we demonstrated that a more interconnected circuit can enhance the model resilience. The work in [12] provides an overview of the vulnerabilities and potential robustness of QML against adversarial attacks, emphasizing the intrinsic properties of Hilbert space dimensions and their implications for adversarial robustness. Building upon this work, our study explores the impact of Hilbert space expressibility and entanglement strength properties from the circuit architecture perspective on the QuNN resilience.
@@ I-A Limitations of State-Of-The-Art and Main Contributions
  • We design the AdvQuNN methodology to test different types of adversarial attacks on QuNNs and classical DNN models, varying the quantum circuit type. (Section III)
  • We compare the robustness of QuNNs to classical CNNs against various types of adversarial attacks. (Section IV-B)
  • We evaluate the impact of the Ansatz architecture on the QuNN robustness (Section IV-C)
Figure 1: Overview of our novel contributions in this work.
II Background and Related Work
II-A Classical Convolutional Neural Networks
Convolutional Neural Networks (CNNs) [13, 14], particularly effective in visual tasks [15] like image classification [16], object detection [17], and semantic segmentation [18], rely on convolutional layers to extract spatial features. These layers apply filters to input images, producing feature maps that represent detected features. The convolution operation is mathematically defined as $F(i,j)=(X*K)(i,j)=\sum_{m}\sum_{n}X(m,n)K(i-m,j-n)$, where $F$ is the feature map, $X$ is the input image, and $K$ is the kernel. Subsequent layers combine these features into more complex patterns, following $F^{l}=\sigma\left(\sum_{k=1}^{K}F_{k}^{l-1}*K_{k}^{l}+b_{k}^{l}\right)$, with $\sigma$ as the activation function.
II-B Quanvolutional Neural Networks
Quanvolutional Neural Networks (QuNNs) [3] represent an intersection between the domains of quantum computing and deep learning that provides an advanced approach to process and understand data. QuNN models substitute the conventional convolutional filter in the CNN with a quantum filter known as the quanvolutional filter. Like classical convolutional layers, the quanvolutional layer applies an operation analogous to an $m\times m$ filter, but it does so by executing an $m^{2}$-qubit quantum circuit (if we use direct pixel encoding as in this work) to produce a downsampled output image. The quanvolutional layer includes three components: data encoding, a parameterized unitary $U(\theta)$, and measurement $M$.
Data encoding. The input data to the quantum layer, consisting of $n$ features, is encoded into an $n$-qubit Hilbert space using quantum data encoding techniques [19]. In this work, we used the angle encoding method where each classical data feature maps to a rotation angle, thus preserving data structure in the quantum circuit. The quantum state of an angle encoded input data $x=[x_{0},x_{1},\dots,x_{n}]^{\intercal}\in\mathbb{R}^{n}$ is represented as
$|\psi_{x}\rangle=\bigotimes_{i}^{n}R\left(\mathbf{\phi}_{i}\right)\left|0^{n}\right\rangle\vspace{-5pt}$ (1)
@@ II-B Quanvolutional Neural Networks
Parameterized unitary $U(\theta)$. A unitary operator $U$, represented by a $2^{n}\times 2^{n}$ matrix, configures an $n$-qubit quantum circuit using single-qubit gates (e.g., ${R_{x},R_{y},R_{z},H,\dots}$) and two-qubit gates (e.g., ${CNOT,CR_{x},CR_{y},CR_{z},\dots}$). These gates enable the creation of various unitary operations $U_{i}$, each forming a distinct quanvolutional filter with specific gate sequences and entanglement patterns. Parameters for these filters can be iteratively updated with classical layer parameters or initialized randomly. In this study, our methodology focuses on the impact of optimizing circuit configuration rather than its parameters, hence, we opt for random initialization. This also will allow us to reduce computational demands and simplify the quantum circuit configuration process. The selected parametric circuit $U(\theta)$ is applied to the initial state as follows:
$|\psi_{x}(\theta)\rangle=U(\theta)|\psi_{x}\rangle$ (2)
@@
$\langle M\rangle=\langle\psi_{x}(\theta)|M|\psi_{x}(\theta)\rangle=\frac{1}{S}\sum_{i}p_{i}\lambda_{i}$ (3)
@@
Figure 2: Quanvolutional Neural Network architecture. A 2 by 2 block of the input image of digit $2$ is encoded through a of rotation gates $R_{y}$, producing the state $|\psi_{x}\rangle$. This state is quanvolved with a parameterized unitary gate $U(\theta)$. Subsequently, we measure the state $|\psi_{x}(\theta)\rangle$ and compute the expectation value across all qubits, and assign values to four channels, forming a feature map. Finally, a sequence of fully connected classical layers then processes this map for classification.
Figure 2 shows an example of a QuNN architecture applied to the MNIST dataset. A $2\times 2$ square of our input image is first encoded with $R_{y}$ gates into the $2\times 2$ quanvolutional kernel of $4$ qubits. This data then undergoes transformation by a unitary $U(\theta)$. Finally, we measure the expectation value of the Pauli-$Z$ observable for each qubit. Each expectation value contributes to a separate channel in the resulting pixel. Therefore, with four qubits, we obtain four channels. These channels are subsequently flattened and input into a classical dense layer, which determines the likelihood of each digit.
II-C Adversarial Attacks
@@
In Quantum Machine Learning (QML), particularly with quanvolutional neural networks, the issue of adversarial attacks presents a dual challenge of vulnerability and potential robustness. The works in [12] and [20] reveal that these networks, whether they process classical or quantum input data, share vulnerabilities with classical networks, while also presenting a trade-off between algorithmic security and quantum advantages. Conversely, studies in [21] and [22] highlight an intrinsic robustness of quantum models to such attacks. The works in [5][23][24], and [25] further confirm the superior performance of QuNNs over classical networks in terms of accuracy and resilience against adversarial attacks both empirically and theoretically. Our study focuses on investigating a novel approach to evaluate and enhance the robustness of QuNNs. The attack algorithms used in this work, FGSM, PGD, and MIM, are shown in Algorithms 12, and 3, respectively.
Algorithm 1 FGSM Attack
1:Model $model$, Data $X$ of size $M$, True Labels $y$, Perturbation size $\epsilon$
2:Adversarial examples $\hat{X}$
3:function FGSM_Attack($model,X,y,\epsilon$)
4:     $gradients\leftarrow$ Compute_Gradient($model,X,y$)
5:     $sign\_gradients\leftarrow\text{sign}(gradients)$
6:     $\hat{X}\leftarrow X+\epsilon\cdot sign\_gradients$
7:     $\hat{X}\leftarrow\text{clip}(\hat{X},0,1)$
8:     return $\hat{X}$ of size $M$
9:end function
10:function Compute_Gradient($model,X,y$)
11:     with GradientTape() as tape:
12:         tape.watch($X$)
13:         prediction $\leftarrow$ model($X$)
14:         loss $\leftarrow$ sparse_categorical_crossentropy($y$, prediction)
15:     gradients $\leftarrow$ tape.gradient(loss, $X$)
16:     return $gradients$
17:end function
Algorithm 2 PGD Attack
1:Model $model$, Data $X$ of size $M$, True Labels $y$, Perturbation size $\epsilon$, Iterations $k$, Step size $\alpha$
2:Adversarial examples $\hat{X}$
3:function PGD_Attack($model,X,y,\epsilon,k,\alpha$)
4:     $\hat{X}\leftarrow X$
5:     for $i=1$ to $k$ do
6:         FGSM_Attack($model,\hat{X},y,\epsilon$)
7:         $\hat{X}\leftarrow\text{clip}(\hat{X},X-\epsilon,X+\epsilon)$
8:         $\hat{X}\leftarrow\text{clip}(\hat{X},0,1)$
9:     end for
10:     return $\hat{X}$ of size $M$
11:end function
Algorithm 3 MIM Attack
1:Model $model$, Data $X$ of size $M$, True Labels $y$, Perturbation size $\epsilon$, Iterations $k$, Step size $\alpha$, Decay factor $\mu$
2:Adversarial examples $\hat{X}$
3:function MIM_Attack($model,X,y,\epsilon,k,\alpha,\mu$)
4:     $\hat{X}\leftarrow X$
5:     $g\leftarrow 0$triangleright Initialize the momentum term
6:     for $i=1$ to $k$ do
7:         $gradients\leftarrow$ Compute_Gradient($model,\hat{X},y$)
8:         $g\leftarrow\mu\cdot g+\frac{gradients}{\|gradients\|_{1}}$triangleright Update momentum term
9:         $sign\_gradients\leftarrow\text{sign}(g)$
10:         $\hat{X}\leftarrow\hat{X}+\alpha\cdot sign\_gradients$
11:         $\hat{X}\leftarrow\text{clip}(\hat{X},X-\epsilon,X+\epsilon)$
12:         $\hat{X}\leftarrow\text{clip}(\hat{X},0,1)$
13:     end for
14:     return $\hat{X}$ of size $M$
15:end function
III AdvQuNN Methodology
@@
In our methodology, we focus on the adaptability of QuNNs by customizing the quanvolutional filter $U(\theta)$, which is crucial for adapting the parametrized quantum circuit, or Ansatz. This versatility led to the development of AdvQuNN, a methodology aimed at evaluating the robustness of QuNNs, paving the way for enhanced methods to robustify QuNNs against various adversarial attacks by fine-tuning components such as gate selection and entanglement settings.
@@
Unlike prior research that used random Ansatz architectures, we evaluated various Ansatz architectures and found that the expressibility and entanglement strength of the quantum circuit, influenced by rotation gates and entanglement configurations [26], impact the feature extraction capability of QuNNs, hence its robustness. As quantum entanglement has an important role in the decision boundary formation of the quanvolved data, we explored different entanglement patterns and identified the five most robust architectures for our dataset (Figure 3), including no entanglement and random architectures [3], as well as architectures that follow the popular nearest-neighbor and all-to-all entanglement topologies. Additionally, the circuit gate selection is inspired by data encoding technique proved effective for data separation in Quantum Support Vector Machine (QSVM) [27], specifically the $ZZ$ feature map that employs $ZZ$ entanglement.
Figure 3: A set of Ansatz architectures considered in the study. Each is labeled as (a) No entanglement, (b) ZZ full entanglement, (c) ZZ linear entanglement, (d) ZZ star entanglement, and (e) Random architecture. The $Rot(a,b,c)$ is a set of three rotation gates $R_{z}(a)R_{y}(b)R_{z}(c)$. The gates $R_{x}$, $R_{y}$, and $R_{z}$ are parameterized.
@@
The $ZZ$ entanglement pattern, involving a controlled phase shift $e^{-i\theta}$ between two qubits. This entanglement is created using the unitary transformation $U_{ZZ}(\theta)=e^{-i\theta Z\otimes Z}$, where $Z$ is the Pauli-$Z$ matrix and $\theta$ is the angle of rotation, as shown in red box Figure 3b.
@@
C) $ZZ$ linear entanglement architecture (Figure 3c). In this architecture, each qubit is entangled linearly with its nearest neighbor, forming a chain of entangled qubits with just $n+1$ operations for $n$ qubits. Its corresponding unitary operator is:
$U_{zzlinear}=U^{\mathrm{linear-ent}}\left(\bigotimes_{0}^{n}Rot\left(a,b,c\right)\right)\vspace{-10pt}$ (5)
$U^{\text{linear-ent}}:=\prod_{q=1}^{n-1}\mathbf{E}_{q,q+1}\;\;\;\text{where}\;\;\;\mathbf{E}_{q,q+1}=e^{-i\theta Z_{q}Z_{q+1}}$ (6)
@@
$U_{zzfull}=U^{\mathrm{full-ent}}\left(\bigotimes_{0}^{n}Rot\left(a,b,c\right)\right)$ (7)
$U^{\text{full-ent }}:=\prod_{q=1}^{n-1}\prod_{k=q+1}^{n}\mathbf{E}_{q,k}\;\;\;\text{where}\;\;\;\mathbf{E}_{q,k}=e^{-i\theta Z_{q}Z_{k}}$ (8)
@@
$U_{zzstar}=U^{\mathrm{star-ent}}\left(\bigotimes_{0}^{n}Rot\left(a,b,c\right)\right)\vspace{-10pt}$ (9)
$U^{\text{star-ent}}:=\prod_{q=1}^{n-1}\mathbf{E}_{1,q+1}\;\;\;\text{where}\;\;\;\mathbf{E}_{1,q+1}=e^{-i\theta Z_{1}Z_{q+1}}\vspace{-5pt}$ (10)
IV Results and Discussion
IV-A Experimental Setup